A stationary store has decided to accept a large shipment of ball-point pens if an inspection of 17 randomly selected pens yields no more than two defective pens.
(a) Find the probability that this shipment is accepted if 5% of the total shipment is defective. (Use 3 decimal places.)

(b) Find the probability that this shipment is not accepted if 15% of the total shipment is defective. (Use 3 decimal places.)

Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

Probability of 0 defective is .95^17=0.4181
probability of 1 defective is 17 (17 ways) .95^16*.05=0.3741
for 2 it is 17C2*,95^15*.05^2=0.1575
Those add to 0.9497 or 0.950.
For the next one, use .85 and .15
.85^17=0.0631
17*.85^16*.15=0.1893
136*.85^15*.15^2=0.2673
That sum is 0.5197

Answer by Alan3354(69443)

A stationary store …
======================
A store that doesn't move.